OpenStudy (anonymous):

Differential Equations: Given that y1(t) = t^4 is a known solution of the second order linear differential equation t^2y'' - 4ty' + 4y = 0, t > 0, find the general solution of the equation.

OpenStudy (anonymous):

let y=x^m, y'=mx^(m-1) y"=m(m-1)x^(m-2) plug in the values in the eqn, m^2-m-4m+4=0 (m-4)(m-1)=0 m=1,4 y=Ax+Bx^4